For 2000 patients, blood-clotting time was normally distributed with a mean of 8 seconds and a standard deviation of 3 seconds. What percent had blood-clotting times between 5 and 11 seconds?
Thanks

Found 2 solutions by Nate, venugopalramana:
Answer by Nate(3500)

Use normal distributation!
--- 2% -- 14% - 34% - 34% - 14% --- 2% --
------- 2 ---- 5 ---- 8 ---- 11 ---- 14 ----------
What percent had blood-clotting times between 5 and 11 seconds?
34% + 34% = 68%

Answer by venugopalramana(3286)

For 2000 patients, blood-clotting time was normally distributed with a mean of 8 seconds and a standard deviation of 3 seconds. What percent had blood-clotting times between 5 and 11 seconds?
MEAN=8
S.D=3
N=2000
S.E.=S.D/SQRT(N)=3/SQRT(2000)=0.06708
VARIATION FROM MEAN =8-3=5 AND 8+3=11
WE FIND 3=0.06708*K
K=44.7
THAT IS WE WANT % LIMITS FOR 44.7SIGMA...IT IS ALMOST 100%
SO ALMOST 100% OF PATIENTS WILL HAVE CLOTTING TIME OF 5 TO 11 SECS.